OpenStudy (anonymous):

I need help with this: -3x-9=-12

OpenStudy (anonymous):

Add 9 to both sides, you get -3x=-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then, divide -3 to both sides

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

ok, ty

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

-3x=-12+9 -3x=-3 -3x/-3=-3/-3 =1

OpenStudy (anonymous):

you're welcome